How would I solve this problem?
Someone invested $1200 at a certain interest rate and $2200 at a rate that is 3% higher. The person's annual earnings were 253. Find the two interest rates.
Thank you!

Answer by stanbon(75887)

Someone invested $1200 at a certain interest rate and $2200 at a rate that is 3% higher. The person's annual earnings were 253. Find the two interest rates.
---
Let the certain rate be "x".
-----
Equation:
interest + interest = interest
------
1200x + 2200(x+.03) = 253
1200x + 2200x + 66 = 253
3400x = 287
x = 0.0844 or 8.44%
x + 0.03 = 0.1144 or 11.44%
=======================
Cheers,
Stan H.
===============